OpenStudy (legends):

Can you simplify 38/64

OpenStudy (legends):

Can you help @Starbucks

OpenStudy (anonymous):

its 19/32

OpenStudy (legends):

How?

OpenStudy (anonymous):

divide both by 2 and because 19 is a prime number thats how its simplified

OpenStudy (yanasidlinskiy):

Find a number that would go evenly into both of the numbers and just divide right out.

OpenStudy (legends):

Thanks

OpenStudy (yanasidlinskiy):

That's all it takes to do.

OpenStudy (anonymous):

Fraction: 38 / 64 = (2 * 19) / 26 = (38 : 2) / (64 : 2) = 19 / 32

OpenStudy (anonymous):

your welcome (: